### v2

@@ Acknowledgements
The authors thank Scott Aaronson and Or Sattath for insightful discussions and comments on the manuscript.
@@ Appendix A Omitted Proofs
    Theorem 9 of [39], applied to square matrices. The proof is based on lower bounds for distributed Fourier sampling.
@@
    Suppose Alice and the Bobs manage to solve $\mathrm{MLRS}_{N,m}$ with inaccuracy $\eta$. This means that Bob produces a sample from a distribution that is at most $\eta$ in TV from each of his target distributions $\mathcal{P}_{j}$. From the definition of $x$ and the $B^{(j)}$, $\mathcal{P}_{j}$ is be a delta function at $r_{j}$. This means that with probability at least $1-2\eta$, Bob recovers the $\log(N/m)$ bits of $r_{j}$ by performing a computational basis measurement. It follows that Alice’s message to Bob is a random-access encoding of $m\log(N/m)$ bits. From known lower bounds on the number of qubits needed for random access coding [41], if $22\eta<1/$ 2, Alice must send at least $\Omega(m\log(N/m))$ qubits to the Bobs.

### v3

@@ Acknowledgements
The authors thank Scott Aaronson, Dmytro Gavinsky, and Or Sattath for insightful discussions and comments on the manuscript. S.J. was supported by Scott Aaronson’s CIQC grant.
@@ Appendix A Omitted Proofs
    Theorem 9 of [MS24], applied to square matrices. The proof is based on lower bounds for distributed Fourier sampling.
@@
    Suppose Alice and the Bobs manage to solve $\mathrm{MLRS}_{N,m}$ with inaccuracy $\eta$. This means that Bob produces a sample from a distribution that is at most $\eta$ in TV from each of his target distributions $\mathcal{P}_{j}$. From the definition of $x$ and the $B^{(j)}$, $\mathcal{P}_{j}$ is be a delta function at $r_{j}$. This means that with probability at least $1-2\eta$, Bob recovers the $\log(N/m)$ bits of $r_{j}$ by performing a computational basis measurement. It follows that Alice’s message to Bob is a random-access encoding of $m\log(N/m)$ bits. From known lower bounds on the number of qubits needed for random access coding [Nay99], if $22\eta<1/$ 2, Alice must send at least $\Omega(m\log(N/m))$ qubits to the Bobs.
@@
Appendix B Discussion
Being a preliminary investigation into the possibility of using quantum networks in this manner, our results do not immediately apply to problems with clear economic value. If this were the case, it could enable novel types of data markets and incentive structures for the production of data. It is worth noting however that our results for the linear regression sampling problem apply also to a related problem in which Bob obtains a state that encodes the solution to a linear system rather than a classical sample. Such states are known to be strictly more powerful resources than classical samples [ATS03], and could potentially be useful in learning tasks such as updating the value of a linear estimator with new data (which is typically achieved with the recursive least squares algorithm).
The form of the quantum communication lower bound that indicates the rival behavior of quantum data is reminiscent of a direct sum theorem. Direct sum theorems demonstrate that the complexity of solving $m$ independent instances of certain problems scales linearly with $m$. They have been studied extensively in both the classical [Sha03, BRWY13, LSS08] and quantum [She11, JK21] setting. These results are not directly applicable since in our setting the inputs to Alice are not independent. Thus, this work motivates an asymmetric direct sum result for classes of communication relations.
In restricting access to data that is used for computation, the setting we consider bears some resemblance to that of differential privacy [DR14, AR19]. In differential privacy, a query is promised not to reveal too much about individual datapoints. This is typically achieved classically by adding noise to data, while we achieve a similar capability in spirit by using a noiseless encoding into quantum states.